# The merger of spinning, accreting supermassive black hole binaries

### v1

@@ Abstract
Introduction—A central goal in multimessenger astronomy is to connect gravitational wave (GW) signals from supermassive binary black hole (SMBBH) systems with their electromagnetic (EM) counterparts [1, 2]. Photon observations are critical for identifying and characterizing these systems all the way from inspiral to merger and its aftermath [3, 4]. For example, such signals may enable the detection of SMBBHs before the GWs can be detected, alert us to recent mergers, and allow us to track their orbital evolution [5, 6]. The unique structure of the accreting gas—often featuring a circumbinary disk (CBD) with a central gap and individual “minidisks”—is expected to generate distinct spectral features and periodic light curves that distinguish these systems from single accreting supermassive black holes, but there is considerable controversy over just what these features may be [7, 8, 9, 10, 11, 4, 12, 13].
Using our recently developed computational tools for simultaneous solution of the MHD and Einstein field equations [14], we have found significant new results about dynamics shortly before, during, and after the merger of a SMBBH with spinning black holes and their implications for photon emission and jet launching. Here we focus on the case in which both spin axes are parallel to the angular momentum of the binary. This simplest spin configuration is favored by evolutionary models [15, 16].
Prior to merger, both black holes can support jets. Remarkably, we find a strong linear correlation between the Poynting luminosity of the jets and the photon luminosity from the gas. Moreover, during the first portion of the inspiral we simulate, when the SMBBH separation is $\sim\!15\text{--}20\,M$} (the gravitational radius $r_{g}\equiv GM/c^{2}=M$} for G=c=1{G=c=1}, so both length and time are in units of $M$), both luminosities are more tightly correlated with the gas mass in which the black holes are immersed than with the mass accretion rate onto the black holes. Over the course of the inspiral, both luminosities decrease gradually by factors of several, but neither disappears; at the moment of merger, both jump back to roughly their original level. Such distinct lightcurves might be used to identify SMBBH mergers even without direct detection of GWs.
@@
Principal results—As the inspiral begins, minidisks quickly form around each black hole (top-left panel of Figure 1). As in [14], soon thereafter each minidisk begins to alternate between a “disk-like” state, with maximum mass and minimum accretion rate onto its black hole, and a “stream-like” state, with opposite properties [22]. However, the initial time-averaged minidisks’ masses are $4\text{--}5\times$} larger than in the nonspinning configuration of [14], mostly due to the smaller ISCO of a spinning black hole. The minidisks also retain their mass for longer: the mass of one takes a time $\sim\!4300\,M$} to fall to a value matching the initial minidisk mass in the nonspinning run, and the other does not fall to that level for another $\sim\!3000\,M$}. This extended lifetime of the minidisks is a consequence of the “hangup” effect [23], in which repulsive spin-orbit coupling both lengthens the inspiral and holds the constituent gas at larger radii before the merger [21].
@@
The tight link between $L_{\text{photons}}$} and $L_{\text{Poynting}}$} can be understood if the gas is in thermal balance with a cooling time $t_{\text{cool}}$} independent of the heating rate, and the magnetic flux on the black holes is confined by the thermal pressure of surrounding gas rather than the ram pressure of accreting mass. Thermal balance should be the case here because the cooling time is comparable to the orbital dynamical time and is independent of temperature, and the inflow time ($M_{\text{gas}}/\dot{M}$}) is generally $\gtrsim 10\times$} longer. The Poynting luminosity $L_{\text{Poynting}}\sim B^{2}M^{2}f\!\left(a/M\right)$} , where $B$ is the poloidal magnetic field on the horizon, the surface area of the black hole is $\sim\!M^{2}$}, and $f\!\left(a/M\right)$} describes how the Poynting luminosity depends on the spin parameter a/M{a/M} [33]. If the magnetic pressure $\sim\!B^{2}$} is matched by the gas pressure $P$, and the gas is in thermal balance, then $P\sim\left(M_{\text{gas}}/r^{3}\right)\left(kT/m_{p}\right)\propto L_{\text{photons}}t_{\text{cool}}/r^{3}$}. It immediately follows that $L_{\text{Poynting}}\propto L_{\text{photons}}$} .

### v3

@@ Abstract
Introduction—A central goal in multimessenger astronomy is to connect gravitational wave (GW) signals from supermassive binary black hole (SMBBH) systems with their electromagnetic (EM) counterparts [1, 2]. Photon observations are critical for identifying and characterizing these systems all the way from inspiral to merger and its aftermath [3, 4]. For example, such signals may enable the detection of SMBBHs before the GWs can be detected, alert us to recent mergers, and allow us to track their orbital evolution [5, 6]. The unique structure of the accreting gas—often featuring a circumbinary disk (CBD) with a central gap and individual “minidisks”—is expected to generate distinct spectral features and periodic light curves that distinguish these systems from single accreting supermassive black holes, but there is considerable controversy over just what these features may be (e.g., [4] for a review).
Some have argued that photon emission should drop sharply as the binary shrinks and not revive until well after merger [7, 8, 9, 10, 11, 12] while others predict that it might drop by a factor $\sim O\left(1\right)$} and then revive quickly after merger [13, 14, 15, 16, 17, 18, 19, 20, 21]. Many have suggested ways the luminosity and jet Poynting flux should generically be modulated periodically [13, 14, 15, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], while some have argued that modulations should be weak or indistinct throughout the inspiral [15] or as merger is approached [17, 30]. Analytic estimates [5, 24] and smoothed particle hydrodynamics (SPH) simulations [22, 10] predict a significant “notch” in the thermal spectrum well before merger, while others have argued this feature does not appear at all [10] and GRMHD simulations of late inspiral indicate the notch may disappear near merger [32, 33, 34, 21]. To date, none of these signals has been observed.
Using our recently developed computational tools for simultaneous solution of the magnetohydrodynamics (MHD) and Einstein field equations [21], we have found significant new results about dynamics shortly before, during, and after the merger of a SMBBH with spinning black holes, and these results have strong implications for photon emission and jet launching. Here we focus on the case in which both spin axes are parallel to the angular momentum of the binary. This simplest spin configuration is favored by evolutionary models [35, 36].
Prior to merger, both black holes can support jets. Remarkably, we find a strong linear correlation between the Poynting luminosity of the jets and the photon luminosity from the gas. Moreover, during the first portion of the inspiral we simulate, when the SMBBH separation is $\sim\!15\text{--}20\,M$} (the gravitational radius $r_{g}\equiv GM/c^{2}=M$} and the gravitational timescale $t_{g}\equiv r_{g}/c=M$} for G=c=1{G=c=1}, so both length and time are in units of $M$), both luminosities are more tightly correlated with the gas mass in which the black holes are immersed than with the mass accretion rate onto the black holes. Over the course of the inspiral, both luminosities decrease gradually by factors of several, but neither disappears; at the moment of merger, both jump back to roughly their original level. Such distinct lightcurves might be used to identify SMBBH mergers even without direct detection of GWs.
@@
Principal results—As the inspiral begins, minidisks quickly form around each black hole (top-left panel of Figure 1). As in [21], soon thereafter each minidisk begins to alternate between a “disk-like” state, with maximum mass and minimum accretion rate onto its black hole, and a “stream-like” state, with opposite properties [20]. However, the initial time-averaged minidisks’ masses are $4\text{--}5\times$} larger than in the nonspinning configuration of [21], mostly due to the smaller radius of the innermost stable circular orbit (ISCO) of a spinning black hole. The minidisks also retain their mass for longer: the mass of one takes a time $\sim\!4300\,M$} to fall to a value matching the initial minidisk mass in the nonspinning run, and the other does not fall to that level for another $\sim\!3000\,M$}. This extended lifetime of the minidisks is a consequence of the “hangup” effect [42], in which repulsive spin-orbit coupling both lengthens the inspiral and holds the constituent gas at larger radii before the merger [41].
@@
The tight link between $L_{\text{photons}}$} and $L_{\text{Poynting}}$} can be understood if the gas is in thermal balance with a cooling time $t_{\text{cool}}$} independent of the heating rate, and the magnetic flux on the black holes is confined by the thermal pressure of surrounding gas rather than the ram pressure of accreting mass. The surrounding gas should be in thermal balance because, independent of temperature, the cooling time is comparable to the orbital dynamical time, and the inflow time ($M_{\text{gas}}/\dot{M}$}) is generally $\gtrsim 10\times$} longer. The Poynting luminosity $L_{\text{Poynting}}\!\sim\!B^{2}M^{2}f\!\left(a/M\right)$} , where $B$ is the poloidal magnetic field on the horizon, the surface area of the black hole is $\sim\!M^{2}$}, and $f\!\left(a/M\right)$} describes how the Poynting luminosity depends on the spin parameter a/M{a/M} [52]. Thermal balance in the gas implies that the gas pressure $P\!\sim\!L_{\text{photons}}t_{\text{cool}}/r^{3}$} because the cooling rate is exactly the photon luminosity; if the magnetic pressure $B^{2}$ is matched by $P$, it immediately follows that $L_{\text{Poynting}}\propto L_{\text{photons}}$} .